OpenStudy (anonymous):

Solve in scientific notation. ( 9 x 10 ^5) (6 x 10 ^-7)

OpenStudy (anonymous):

( 9 x 10 ^5) (6 x 10 ^-7) 900000(0.0000006)=.54 5.4 x 10^-1

OpenStudy (anonymous):

thank you!

OpenStudy (unklerhaukus):

\[( 9 \times 10 ^5) (6 \times 10 ^{-7})\]\[=9\times6\times 10^{5-7}\]